OpenStudy (anonymous):

Which of the following is the sketch of y2 – 8y – 12x – 8 = 0?

OpenStudy (campbell_st):

it looks like a parabola which is concave right...  \[y^2 - 8y = 12x + 8\] complete the square \[y^2 - 8y + 16 = 12x +8 + 16\] which can be factorised to \[(y -4)^2 = 12(x +2)\] the general form is \[(y -k)^2 = 4a(x -h)\] the vertex is (h, k) and focal length is a hope this helps